OpenStudy (anonymous):

I will medal! Please help! Solve 8x2 + 7x - 15 = - 7.

OpenStudy (anonymous):

Please help.

Nnesha (nnesha):

set it equal to zero move -7 to the left side

Nnesha (nnesha):

add 7 both sides \[ \huge\rm 8x^2 + 7x - 15\color{reD}{+7} = - 7\color{reD}{+7}\]\[ \large\rm 8x^2 + 7x \color{reD}{- 8} = 0\] now you can use quadratic formula to solve for x \[\huge\rm \frac{ {- }b \pm \sqrt{b^2 -4ac} }{ {2a} }\] quadratic formula where abc values are \[\large\rm Ax^2 +Bx+C=0\]

Nnesha (nnesha):

\[ \large\rm \color{red}{8}x^2 +\color{Red}{ 7}x \color{reD}{- 8} = 0\]\[\large\rm \color{blue}{A}x^2 +\color{blue}{B}x+\color{blue}{C}=0\]

Nnesha (nnesha):

or ofc you can factor it